djmatthies12:

Write the equation of the line that passes through (3, 4) and (2, −1) in slope-intercept form. (2 points) y = 3x − 7 y = 3x − 5 y = 5x − 11 y = 5x − 9

forgetmylife:

Step one : In order to find the equation of the line we can use the slope intercept form. The slope-intercept form of a straight line is used to find the equation of a line. For the slope-intercept formula, we have to know the slope of the line and the intercept cut by the line with the y-axis. y = mx + b Where m is the slope and b is the y-intercept We know that, m = (y2 - y1)/ (x2 - x1)

forgetmylife:

Substituting the values we get, m = (-1 - 4)/(2 - 3)

forgetmylife:

So m=-5/-1 Dividing two negative numbers becomes positive so m=5

forgetmylife:

Step 2 to find b we’ll use the point (2,1) The formula used is y=mx+c

forgetmylife:

Which in that case -1=5(2)+b b=-10-1 (adding two negatives will remain negative) b=-11

forgetmylife:

So now what would your answer be?